OpenStudy (anonymous):

f(x) = ln x, [1, 8] Using mean value theorem, find all numbers c that satisfy the conclusion of the Mean Value Theorem.

myininaya (myininaya):

\[f'(c)=\frac{f(b)-f(a)}{b-a}\]

OpenStudy (anonymous):

I found my c to be 3.366 but the answer is wrong

myininaya (myininaya):

a=1 and b=8

OpenStudy (anonymous):

yes, i know how to do it..but got it wrong.. don't know why

myininaya (myininaya):

Does it want an approximation?

OpenStudy (anonymous):

it didn't say..so i'm thinking should the answer be an exact number

OpenStudy (anonymous):

(Enter your answers as a comma-separated list. If it does not satisfy the hypotheses, enter DNE).

OpenStudy (anonymous):

but then i think there is only 1 number c

myininaya (myininaya):

Well the number you put above is not the exact answer

myininaya (myininaya):

That is call an appoximation

OpenStudy (anonymous):

how do get an exact number?

myininaya (myininaya):

So I assume you did the setup right because you got the approximation to the correct answer

myininaya (myininaya):

This is the setup you chose? \[\frac{1}{c}=\frac{\ln(8)-\ln(1)}{8-1}\]

OpenStudy (anonymous):

yes

myininaya (myininaya):

We know ln(1) =? and 8-1=? So can you tell me what 1/c= Don't use your calculator

OpenStudy (anonymous):

and then f'(c) = 1/c then solve for c

OpenStudy (anonymous):

ln 1 = 0 8-1= 7

myininaya (myininaya):

Ok we have \[\frac{1}{c}=\frac{\ln(8)-0}{7}\] Do you know how to solve this for c?

OpenStudy (anonymous):

ln 8/ 7 = 1/c c= 7/ln8

myininaya (myininaya):

right

OpenStudy (anonymous):

ok thanks so much!